What mass of Na2S2O3 is needed to react with 0.12 g of Cl2?
1
Expert's answer
2020-12-01T09:08:19-0500

Solution:

The balanced chemical equation:

Na2S2O3(aq) + 4Cl2(g) + 5H2O(aq) --> 2NaHSO4(aq) + 8HCl(aq)

According to the equation: n(Na2S2O3) = n(Cl2)/4


Moles of Cl2 = n(Cl2) = Mass of Cl2 / Molar mass of Cl2

The molar mass of Cl2 is 70.9 g mol-1.

n(Cl2) = 0.12 g / 70.9 g mol-1 = 0.00169 mol = 0.0017 mol


n(Na2S2O3) = n(Cl2)/4

n(Na2S2O3) = 0.0017 mol / 4 = 0.000425 mol


Moles of Na2S2O3 = n(Na2S2O3) = Mass of Na2S2O3 / Molar mass of Na2S2O3

The molar mass of Na2S2O3 is 158.11 g mol-1.

Hence,

Mass of Na2S2O3 = n(Na2S2O3) × M(Na2S2O3)

Mass of Na2S2O3 = 0.000425 mol × 158.11 g mol-1 = 0.067 g

Mass of Na2S2O3 is 0.067 grams.


Answer: 0.067 g of Na2S2O3 is needed to react with 0.12 g of Cl2.